You run a medium-sized farming equipment repair firm in North Dakota. Your busiest season is the late fall through winter months when area farmers need repairs done to their equipment. Because you have limited cash flow, keeping extra people on your payroll is expensive and undesirable. You have determined that each worker must work 120 hours per month and you can repair an average of 42 machines each month. If their labor productivity rate is 0.175, how many workers can you employ at your company?

a. 1.46

b. 1.75

c. 2.0

d. 1.17
Answer #1

Let,
Required number of labours = x

with labor productivity rate of 0.175, x number of labours prepare x*0.175*120 machines per month in 120 hours

So, x*0.175*120 = 42

or, x = 42/(0.175*120) = 2 (Ans)